# MAV: A Multimodal, Multigenre, and Multipurpose Audio-Visual Academic Lecture Dataset

### v1

@@ Appendix C Qualitative Analyses of SSG Benchmark Systems
The comparisons of the benchmark systems in "SlideScript" are shown in Figure 6 and Figure 7. We can find that the scripts generated by "GPT-4+OCR and GPT-4V" are highly qualified and roughly match the references. The results of "LLaMA-2+OCR" are acceptable, although slightly weaker than the previous two. All three can develop detailed descriptions around topics in slides, such as the "NIH Outreach Toolkit" in Figure 6 and "Aging and Migration" in Figure 7. InstructBLIP, on the other hand, fails to find the right theme, as shown in Figure 6 and Figure 7, suggesting that today’s open-source models have difficulty with text in slides.
The comparisons of the benchmark systems in "ScriptSlide" are shown in Figure 9 and Figure 10. The slides generated by GPT-4 are closer to references than those generated by LLaMA-2, and the contents are more accurate and condensed. For example, GPT-4 presents "Drosophila ToLL" in Figure 9 and "Pipeline" in Figure 10, while LLaMA-2 does not. We can conclude that GPT-4 has a better understanding of complex academic knowledge than LLaMA-2.
Figure 8 and Figure 11 show the improvement in the quality of the generated scripts and slides brought by external knowledge. The addition of external knowledge enables AI models to produce necessary details, such as "WiFi CSI" in Figure 8 and "13 long-term users" in Figure 11. It shows that the introduction of external knowledge leads to more comprehensive information captured by AI models.

### v2

@@ Appendix C Qualitative Analyses of SSG Benchmark Systems
The comparisons of the benchmark systems in “SlideScript” are shown in Figure 6 and Figure 7. LLaMA-2 and InstructBLIP below refer to models of size 7B. We can find that the scripts generated by GPT-4+OCR and GPT-4V are highly qualified and roughly match the references. The results of LLaMA-2+OCR are acceptable, although slightly weaker than the previous two. All three can develop detailed descriptions around topics in slides, such as the “NIH Outreach Toolkit” in Figure 6 and “Aging and Migration” in Figure 7. InstructBLIP, on the other hand, fails to find the right theme, as shown in Figure 6 and Figure 7, suggesting that today’s open-source LMM models struggle in handling text in slides.
The comparisons of the benchmark systems in “ScriptSlide” are shown in Figure 9 and Figure 10. The slides generated by GPT-4 are closer to references than those generated by LLaMA-2, and the contents are more accurate and condensed. For example, GPT-4 presents “Drosophila ToLL” in Figure 9 and “Pipeline” in Figure 10, while LLaMA-2 does not. We can conclude that GPT-4 has a better understanding of complex academic knowledge than LLaMA-2.
Figure 8 and Figure 11 show the improvement in the quality of the generated scripts and slides brought by external knowledge. The addition of external knowledge enables AI models to produce necessary details, such as “WiFi CSI” in Figure 8 and “13 long-term users” in Figure 11. It shows that the introduction of external knowledge leads to more comprehensive information captured by AI models.